# A motivic Fundamental Lemma

### v3

@@ 2.13. Orbifold volume
We can assume that ${\mathbf{M}}$ is parametrizing a global quotient $[X/\Gamma]$, where ${\mathbf{X}}$ and $\Gamma$ are as in the beginning of the section. Let $\pi\colon{\mathbf{X}}\to{\mathbf{X}}/\Gamma$ the quotient morphism. Recall the $\mathbf{U}\subseteq{\mathbf{X}}/\Gamma$ is defined as the maximal open subset such that $\pi\colon V=\pi^{-1}(\mathbf{U})\to\mathbf{U}$ is a $\Gamma$-torsor, which amounts to define ${\mathbf{V}}$ as the subset of ${\mathbf{X}}$ where the $\Gamma$-action in free. In this situation, $e$ admits the following more explicit description (see [22, Construction 2.18]). Let $x\in{\mathbf{M}}^{\natural}(K)={\mathbf{M}}(K\mathopen{[\![}t\mathopen{]\!]})\cap\mathbf{U}(K\mathopen{(\!(}t\mathopen{)\!)})$. By definition of $\mathbf{U}$, the point $x$ defines a $\Gamma$-torsor $T$ over $\operatorname{\mathsf{Spec}}(K\mathopen{(\!(}t\mathopen{)\!)})$. Let $\widetilde{T}$ be the normalization of $\operatorname{\mathsf{Spec}}(K\mathopen{[\![}t\mathopen{]\!]})$ in $T$. By the valuative criterion for properness, one gets a $\Gamma$-equivariant map $\widetilde{x}\colon\widetilde{T}\to X$. We have the following diagram:
@@
hence a point in $I_{\hat{\mu}}[X/\Gamma]_{K}$. By [22, Construction 2.18], this point in $I_{\hat{\mu}}M_{K}$ is $e(x)$.
@@
Let us recall the definition of weights from [22, Definition 2.20]. Let $k$ be a field and let $(\chi_{1},\cdots,\chi_{r})\in(\mathbb{Q}/\mathbb{Z})^{r}$. Define $w(\chi_{1},\cdots,\chi_{r})=\sum_{i=1}^{r}c_{i}$, with $c_{i}$ the unique rational representative of $\chi_{i}$ with $0<c_{i}\leq 1$. For any $r$-dimensional $k$-vector space $V$ with an algebraic action $a:\hat{\mu}\to\mu_{N}\to\operatorname{\mathsf{Aut}}(V)$, one has a character decomposition $V_{\bar{k}}\simeq\oplus_{i=1}^{r}V_{\bar{k}}(\chi_{i})$ with $\chi_{i}\in\mathbb{Q}/\mathbb{Z}$ characters of $\widehat{\mu}$ and one defines the weight of $a$ as $w(a)=w(\chi_{1},\cdots,\chi_{r})$. Let ${\mathbf{M}}$ be a definable smooth Deligne-Mumford stack over $k\mathopen{[\![}t\mathopen{]\!]}$ and $(x,\alpha)\in I_{\hat{\mu}}{\mathbf{M}}(K)$. We define $w(x,\alpha)$ as the weight of $\widehat{\mu}$ acting on the tangent space $T_{x}\mathcal{M}_{K}$.
@@
This is the exact analogue for motivic integrals of [22, Theorem 2.21] which deals with $p$-adic integrals. The proof in loc. cit., which ultimately reduces to computations done in [14], adapts without difficulty to the motivic setting, using [34, Proposition 3.2.3] instead of [22, Lemma 4.19]. Note that the apparent discrepancy with [22, Theorem 2.21] (the presence of a denominator in the formula) is an artefact coming from the fact that in [22] $(x,\alpha)$ is a point of $I_{\hat{\mu}}\mathcal{M}$ while here it is a point in the coarse moduli space.
@@ 3. Geometric setup
In this section we introduce the relevant constructions for the formulation of the Geometric Stabilization theorem [37, Théorème 6.4.1]. The main reference is Section 4 of loc. cit.
Throughout this section let $k$ be a field of characteristic 00. Since a first order sentence in the language of rings holds in every pseudo-finite field of characteristic zero if and only it holds for all finite fields of large enough characteristic, see [2], we shall use freely that first order statements proved in [37, 22] for finite fields, sometimes with a lower bound on the characteristic, will hold for any pseudo-finite field containing $k$. We fix $X$ a smooth projective geometrically connected curve over $k$ and $D$ a line bundle of even degree $d$ on $X$. By abuse of notation we also write $D$ for the ${\mathbb{G}}_{m}$-torsor associated with $D$. We will also assume the existence of a rational point $\infty\in X(k)$.
@@ 3.1. Reductive group schemes
We further denote by $T={\mathbb{T}}\times^{\operatorname{\mathrm{Out}}({\mathbb{G}})}\rho$ the induced maximal torus and by $\mathrm{g}$,$\mathrm{t}$,$\mathfrak{g}$ and $\mathfrak{t}$ the Lie algebras of ${\mathbb{G}}$, ${\mathbb{T}}$, $G$ and $T$ respectively. The superscripts $\operatorname{reg}$ and $\operatorname{rs}$ will denote the open dense subschemes of regular and regular semi-simple elements respectively.
@@ 3.2. Higgs bundles
A $G$-Higgs bundle is regular if the morphism $X\rightarrow[\mathfrak{g}_{D}/G]$ factors through $[\mathfrak{g}^{\operatorname{reg}}_{D}/G]$. We denote by ${\mathbb{M}}_{G}^{\operatorname{reg}}$ and ${\mathcal{M}}_{G}^{\operatorname{reg}}$ the corresponding open substacks.
@@ 3.3. The Hitchin fibration
induced by the quotient $[\mathfrak{g}_{D}/G]\to\mathfrak{c}_{D}$ is called the Hitchin fibration.
@@ 3.5. $\pi_{0}({\mathbb{P}}_{G})$ and the anisotropic locus
with $\operatorname{\mathsf{A}}^{\heartsuit}$ the open subset of $\operatorname{\mathsf{A}}$ where the cameral cover is reduced, see [37, 4.5].
@@
${\widetilde{\operatorname{\mathsf{A}}}^{{\mathrm{ani}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widetilde{\operatorname{\mathsf{A}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathfrak{t}^{\operatorname{reg}}_{D,\infty}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{\mathsf{A}}^{{\mathrm{ani}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{\mathsf{A}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{\mathrm{ev}}_{\infty}}$${\mathfrak{c}_{D,\infty}.}$
@@
By [22, Proposition 4.38] the composition of $\phi$ with the projection onto $\pi_{0}(\widetilde{{\mathcal{P}}}_{G})$ gives for every $a\in\widetilde{\operatorname{\mathsf{A}}}^{{\mathrm{ani}}}(k)$ a surjection
@@ 3.6. Coendoscopic groups and twisted inertia
This is a summary of [22, Section 5]. Roughly speaking a coendoscopic group of $G$ is the Langlands-dual to an endoscopic group of $G$.
@@
Coendoscopic groups appear naturally in the study of the twisted inertia stack $I_{\hat{\mu}}{\mathbb{M}}_{G}$ as follows. In [22, Construction 5.6] a natural morphism $\mu_{\mathcal{E}}:[\mathfrak{h}_{{\mathcal{E}},D}/H_{\mathcal{E}}]\to[\mathfrak{g}_{D}/G]$ is constructed, giving rise to compatible morphisms $\mu_{\mathcal{E}}:{\mathbb{M}}_{H_{\mathcal{E}}}\to{\mathbb{M}}_{G}$ and $\nu_{{\mathcal{E}}}:A_{H_{\mathcal{E}}}\to A_{G}$.
Furthermore ${\mathcal{E}}$ induces for any $H_{\mathcal{E}}$-Higgs bundle $F$ a morphism $\hat{\mu}\to\operatorname{\mathsf{Aut}}(F)$, which allows us to lift $\mu_{\mathcal{E}}$ to a morphism to $I_{\hat{\mu}}{\mathbb{M}}_{G}$. In summary we have a commutative diagram
@@
where the vertical arrows are the Hitchin fibrations for $H_{\mathcal{E}}$ and $G$ respectively. Furthermore $\nu_{\mathcal{E}}$ respects the anisotropic locus i.e. $\nu^{-1}_{\mathcal{E}}(A^{{\mathrm{ani}}}_{G})=A^{{\mathrm{ani}}}_{H_{\mathcal{E}}}$.
@@
Now assume further that the fiber of $\rho_{\kappa}$ over $\infty$ has a $k$-rational point $\infty_{\rho_{\kappa}}$. We write $A^{G-\infty}_{H_{\mathcal{E}}}=\nu^{-1}_{\mathcal{E}}(A_{G}^{\infty})$. One can check that $A^{G-\infty}_{H_{\mathcal{E}}}$ is contained in $A^{\infty}_{H_{\mathcal{E}}}$ and we define $\widetilde{A}^{G-\infty}_{H_{\mathcal{E}}}=A^{G-\infty}_{H_{\mathcal{E}}}\times_{A^{\infty}_{H_{\mathcal{E}}}}\widetilde{A}_{H_{{\mathcal{E}}}}$.
@@
The chosen point $\infty_{\rho_{\kappa}}$ induces an isomorphism $\mathfrak{t}_{{\mathcal{E}},\infty}\cong\mathfrak{t}_{\infty}$ which allows us to lift $\nu_{\mathcal{E}}$ to a morphism $\widetilde{\nu}:\widetilde{A}^{G-\infty}_{H_{\mathcal{E}}}\to\widetilde{A}_{G}$.
@@
We further restrict the above constructions to the anisotropic locus, which we indicate by adding a superscript ${\mathrm{ani}}$. For the various exponents * appearing above we will also denote by $\widetilde{{\mathbb{M}}}_{G}^{*}\to\widetilde{A}^{*}_{G}$ and $\widetilde{{\mathbb{M}}}^{*}_{H_{\mathcal{E}}}\to\widetilde{A}^{*}_{H_{\mathcal{E}}}$ the pullbacks of ${\mathbb{M}}_{G}\to A_{G}$ and ${\mathbb{M}}_{H_{\mathcal{E}}}\to A_{H_{\mathcal{E}}}$ to $\widetilde{A}^{*}$ and $\widetilde{A}^{*}_{H_{\mathcal{E}}}$. In particular from (3.6.1) we get the commutative square
[…]
Conversely, if we assume the existence of a $k$-rational point $\infty_{\rho}$ in the fiber of $\rho$ at $\infty$, we get for any anisotropic $G$-Higgs bundle $F$ an inclusion $\operatorname{\mathsf{Aut}}(F)\to{\mathbb{T}}$ [22, Construction 4.36]. Thus any $k$-point in $I_{\hat{\mu}}\widetilde{{\mathbb{M}}}_{G}^{{\mathrm{ani}}}$ defines a morphism $\kappa:\hat{\mu}\to{\mathbb{T}}$ and in fact a coendoscopic datum ${\mathcal{E}}=(\kappa,\rho_{\kappa},\rho_{\kappa}\to\rho)$ [22, Proposition 5.12]. We say that a coendoscopic datum ${\mathcal{E}}$ occurs in $I_{\hat{\mu}}\widetilde{{\mathbb{M}}}_{G}^{{\mathrm{ani}}}(k)$ if it arises from a point in $I_{\hat{\mu}}\widetilde{{\mathbb{M}}}_{G}^{{\mathrm{ani}}}(k)$ via this construction and write $I_{\hat{\mu}}\widetilde{{\mathbb{M}}}_{G}^{{\mathrm{ani}}}(k)_{\mathcal{E}}$ for the component with given occuring endoscopy datum ${\mathcal{E}}$.
Assuming that the fiber of $\rho_{\kappa}$ over $\infty$ has a $k$-rational point, the above constuction defines an equivalence [22, Corollary 5.17]
@@
(3.6.2) $\widetilde{\mu}_{\mathcal{E}}:\widetilde{{\mathbb{M}}}_{H_{\mathcal{E}}}^{G-\infty,{\mathrm{ani}}}(k)\xrightarrow{\>\sim\>}I_{\hat{\mu}}\widetilde{{\mathbb{M}}}_{G}^{{\mathrm{ani}}}(k)_{\mathcal{E}}.$
@@
For our applications we need a variant of this equivalence for the inertia stack of the rigidification $I_{\hat{\mu}}\widetilde{{\mathcal{M}}}_{G}^{{\mathrm{ani}}}$, which will have fewer components indexed by orbits $[{\mathcal{E}}]$ of occuring coendoscopy data under a natural $Z(X,G)$-action [22, Construction 5.20]. Since the coendoscopic groups for two coendoscopy data in the same orbit are isomorphic we will ignore this subtlety and write $I_{\hat{\mu}}\widetilde{{\mathcal{M}}}_{G}^{{\mathrm{ani}}}(k)_{\mathcal{E}}$ for the corresponding component. If we write ~$\widetilde{{\mathcal{M}}}_{H_{\mathcal{E}}}^{G}$ for the $Z(X,G)$-rigidification of $\widetilde{{\mathbb{M}}}_{H_{\mathcal{E}}}^{G-\infty,{\mathrm{ani}}}$, we deduce from (3.6.2) the equivalence [22, Corollary 5.25]
(3.6.3) $\overline{\mu}_{{\mathcal{E}}}:\widetilde{{\mathcal{M}}}_{H_{\mathcal{E}}}^{G}(k)\xrightarrow{\>\sim\>}I_{\hat{\mu}}\widetilde{{\mathcal{M}}}_{G}^{{\mathrm{ani}}}(k)_{\mathcal{E}}.$
@@
Notice also that because of (3.6.1) the equivalences (3.6.2) and (3.6.3) are compatible with the restiriction to individual Hitchin fibers.

### v4

@@ 2.13. Orbifold volume
We can assume that $\mathbf{M}$} is parametrizing a global quotient $X/\Gamma]$, where $\mathbf{X}$} and $\Gamma$ are as in the beginning of the section. Let $\pi\colon{\mathbf{X}}\to{\mathbf{X}}/\Gamma$ the quotient morphism. Recall the $\mathbf{U}\subseteq{\mathbf{X}}/\Gamma$ is defined as the maximal open subset such that $\pi\colon V=\pi^{-1}(\mathbf{U})\to\mathbf{U}$ is a $\Gamma$-torsor, which amounts to define $\mathbf{V}$} as the subset of $\mathbf{X}$} where the $\Gamma$-action in free. In this situation, $e$ admits the following more explicit description (see [23, Construction 2.18]). Let $x\in{\mathbf{M}}^{\natural}(K)={\mathbf{M}}(K\mathopen{[\![}t\mathopen{]\!]})\cap\mathbf{U}(K\mathopen{(\!(}t\mathopen{)\!)})$. By definition of $\mathbf{U}$, the point $x$ defines a $\Gamma$-torsor $T$ over $\operatorname{\mathsf{Spec}}(K\mathopen{(\!(}t\mathopen{)\!)})$. Let $\widetilde{T}$ be the normalization of $\operatorname{\mathsf{Spec}}(K\mathopen{[\![}t\mathopen{]\!]})$ in $T$. By the valuative criterion for properness, one gets a $\Gamma$-equivariant map $\widetilde{x}\colon\widetilde{T}\to X$. We have the following diagram:
@@
hence a point in $I_{\hat{\mu}}[X/\Gamma]_{K}$. By [23, Construction 2.18], this point in $I_{\hat{\mu}}M_{K}$ is $e(x)$.
@@
Let us recall the definition of weights from [23, Definition 2.20]. Let $k$ be a field and let $(\chi_{1},\cdots,\chi_{r})\in(\mathbb{Q}/\mathbb{Z})^{r}$. Define $w(\chi_{1},\cdots,\chi_{r})=\sum_{i=1}^{r}c_{i}$, with $c_{i}$ the unique rational representative of $\chi_{i}$ with $0<c_{i}\leq 1$. For any $r$-dimensional $k$-vector space $V$ with an algebraic action $a:\hat{\mu}\to\mu_{N}\to\operatorname{\mathsf{Aut}}(V)$, one has a character decomposition $V_{\bar{k}}\simeq\oplus_{i=1}^{r}V_{\bar{k}}(\chi_{i})$ with $\chi_{i}\in\mathbb{Q}/\mathbb{Z}$ characters of $\widehat{\mu}$ and one defines the weight of $a$ as $w(a)=w(\chi_{1},\cdots,\chi_{r})$. Let $\mathbf{M}$} be a definable smooth Deligne-Mumford stack over $k\mathopen{[\![}t\mathopen{]\!]}$ and $(x,\alpha)\in I_{\hat{\mu}}{\mathbf{M}}(K)$. We define $w(x,\alpha)$ as the weight of $\widehat{\mu}$ acting on the tangent space $T_{x}\mathcal{M}_{K}$.
@@
This is the exact analogue for motivic integrals of [23, Theorem 2.21] which deals with $p$-adic integrals. The proof in loc. cit., which ultimately reduces to computations done in [14], adapts without difficulty to the motivic setting, using [35, Proposition 3.2.3] instead of [23, Lemma 4.19]. Note that the apparent discrepancy with [23, Theorem 2.21] (the presence of a denominator in the formula) is an artefact coming from the fact that in [23] $(x,\alpha)$ is a point of $I_{\hat{\mu}}\mathcal{M}$ while here it is a point in the coarse moduli space.
@@ 3. Geometric setup
In this section we introduce the relevant constructions for the formulation of the Geometric Stabilization theorem [38, Théorème 6.4.1]. The main reference is Section 4 of loc. cit.
Throughout this section let $k$ be a field of characteristic 0. Since a first order sentence in the language of rings holds in every pseudo-finite field of characteristic zero if and only it holds for all finite fields of large enough characteristic, see [2], we shall use freely that first order statements proved in [38, 23] for finite fields, sometimes with a lower bound on the characteristic, will hold for any pseudo-finite field containing $k$. We fix $X$ a smooth projective geometrically connected curve over $k$ and $D$ a line bundle of even degree $d$ on $X$. By abuse of notation we also write $D$ for the ${\mathbb{G}}_{m}$-torsor associated with $D$. We will also assume the existence of a rational point $\infty\in X(k)$.
@@ 3.1. Reductive group schemes
We further denote by $T={\mathbb{T}}\times^{\operatorname{\mathrm{Out}}({\mathbb{G}})}\rho$ the induced maximal torus and by $\mathrm{g}$$\mathrm{t}$$\mathfrak{g}$ and $\mathfrak{t}$ the Lie algebras of $\mathbb{G}$}, $\mathbb{T}$}, $G$ and $T$ respectively. The superscripts $\operatorname{reg}$ and $\operatorname{rs}$ will denote the open dense subschemes of regular and regular semi-simple elements respectively.
@@ 3.2. Higgs bundles
A $G$-Higgs bundle is regular if the morphism $X\rightarrow[\mathfrak{g}_{D}/G]$ factors through $\mathfrak{g}^{\operatorname{reg}}_{D}/G]$. We denote by ${\mathbb{M}}_{G}^{\operatorname{reg}}$ and ${\mathcal{M}}_{G}^{\operatorname{reg}}$ the corresponding open substacks.
In Section 3.6 below we will also need a slight modification of the notion of a Higgs bundle introduced in [22]. Given a $Z(G)$-gerbe $\beta$ on $X$ we define a $\beta$-twisted $G$-Higgs bundle to be a morphism $X\to[\mathfrak{g}_{D}/G]^{\beta}$, where $\mathfrak{g}_{D}/G]^{\beta}$ is the $\beta$-twist of the $Z(G)$-gerbe $\mathfrak{g}_{D}/G]\to[\mathfrak{g}_{D}/(G/Z(G))]$. We write ${\mathbb{M}}_{G}^{\beta}$ for the corresponding moduli stack.
@@ 3.3. The Hitchin fibration
induced by the quotient $\mathfrak{g}_{D}/G]\to\mathfrak{c}_{D}$ is called the Hitchin fibration.
@@ 3.5. $\pi_{0}({\mathbb{P}}_{G})$ and the anisotropic locus
with $\operatorname{\mathsf{A}}^{\heartsuit}$ the open subset of $\operatorname{\mathsf{A}}$ where the cameral cover is reduced, see [38, 4.5].
@@
[…]
@@
By [23, Proposition 4.38] the composition of $\phi$ with the projection onto $\pi_{0}(\widetilde{{\mathcal{P}}}_{G})$ gives for every $a\in\widetilde{\operatorname{\mathsf{A}}}^{{\mathrm{ani}}}(k)$ a surjection
@@ 3.6. Coendoscopic groups and twisted inertia
This is a summary of [23, Section 5] and [22, Section 2]. Roughly speaking a coendoscopic group of $G$ is the Langlands-dual to an endoscopic group of $G$.
@@
Coendoscopic data appear naturally in the study of the twisted inertia stack $I_{\hat{\mu}}{\mathbb{M}}_{G}$ as follows. Given $\mathcal{E}$} one obtains as in [22, Section 2.3] a $Z(H_{\mathcal{E}})$-gerbe $\beta_{\mathcal{E}}$ and we write $\mathbb{M}$$_{\mathcal{E}}$ for the moduli stack ${\mathbb{M}}_{H_{{\mathcal{E}}}}^{\beta_{\mathcal{E}}}$ of $\beta_{\mathcal{E}}$-twisted $H_{\mathcal{E}}$-Higgs bundles. As in [22, Construction 2.21] we obtain compatible morphisms $\mu_{\mathcal{E}}:{\mathbb{M}}_{\mathcal{E}}\to{\mathbb{M}}_{G}$ and $\nu_{{\mathcal{E}}}:\operatorname{\mathsf{A}}_{\mathcal{E}}=\operatorname{\mathsf{A}}_{H_{\mathcal{E}}}\to\operatorname{\mathsf{A}}_{G}$.
Furthermore $\mathcal{E}$} induces for any $\beta_{\mathcal{E}}$-twisted $H_{\mathcal{E}}$-Higgs bundle $F$ a morphism $\hat{\mu}\to\operatorname{\mathsf{Aut}}(F)$, which allows us to lift $\mu_{\mathcal{E}}$ to a morphism to $I_{\hat{\mu}}{\mathbb{M}}_{G}$. In summary we have a commutative diagram
@@
where the vertical arrows are the Hitchin fibrations. Furthermore $\nu_{\mathcal{E}}$ respects the anisotropic locus i.e. $\nu^{-1}_{\mathcal{E}}(\operatorname{\mathsf{A}}^{{\mathrm{ani}}}_{G})=\operatorname{\mathsf{A}}^{{\mathrm{ani}}}_{H_{\mathcal{E}}}=\operatorname{\mathsf{A}}^{{\mathrm{ani}}}_{{\mathcal{E}}}$.
@@
By [22, Lemma 2.25] the fiber of $\rho_{\kappa}$ over $\infty$ has a $k$-rational point $\infty_{\rho_{\kappa}}$. We write $\operatorname{\mathsf{A}}^{G-\infty}_{{\mathcal{E}}}=\nu^{-1}_{\mathcal{E}}(\operatorname{\mathsf{A}}_{G}^{\infty})$. One can check that $\operatorname{\mathsf{A}}^{G-\infty}_{{\mathcal{E}}}$ is contained in $\operatorname{\mathsf{A}}^{\infty}_{{\mathcal{E}}}$ and we define $\widetilde{\operatorname{\mathsf{A}}}^{G-\infty}_{{\mathcal{E}}}=\operatorname{\mathsf{A}}^{G-\infty}_{{\mathcal{E}}}\times_{\operatorname{\mathsf{A}}^{\infty}_{{\mathcal{E}}}}\widetilde{\operatorname{\mathsf{A}}}_{{{\mathcal{E}}}}$.
@@
The chosen point $\infty_{\rho_{\kappa}}$ induces an isomorphism $\mathfrak{t}_{{\mathcal{E}},\infty}\cong\mathfrak{t}_{\infty}$ which allows us to lift $\nu_{\mathcal{E}}$ to a morphism $\widetilde{\nu}:\widetilde{\operatorname{\mathsf{A}}}^{G-\infty}_{{\mathcal{E}}}\to\widetilde{\operatorname{\mathsf{A}}}_{G}$.
@@
We further restrict the above constructions to the anisotropic locus, which we indicate by adding a superscript $\mathrm{ani}$}. For the various exponents * appearing above we will also denote by $\widetilde{{\mathbb{M}}}_{G}^{*}\to\widetilde{\operatorname{\mathsf{A}}}^{*}_{G}$ and $\widetilde{{\mathbb{M}}}^{*}_{{\mathcal{E}}}\to\widetilde{\operatorname{\mathsf{A}}}^{*}_{{\mathcal{E}}}$ the pullbacks of ${\mathbb{M}}_{G}\to\operatorname{\mathsf{A}}_{G}$ and ${\mathbb{M}}_{{\mathcal{E}}}\to\operatorname{\mathsf{A}}_{{\mathcal{E}}}$ to $\widetilde{\operatorname{\mathsf{A}}}^{*}$ and $\widetilde{\operatorname{\mathsf{A}}}^{*}_{{\mathcal{E}}}$. In particular, from (3.6.1) we get the commutative square
@@
${\widetilde{{\mathbb{M}}}_{{\mathcal{E}}}^{G-\infty,{\mathrm{ani}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widetilde{\mu}_{\mathcal{E}}}$${I_{\hat{\mu}}\widetilde{{\mathbb{M}}}_{G}^{{\mathrm{ani}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widetilde{\operatorname{\mathsf{A}}}_{{\mathcal{E}}}^{G-\infty,{\mathrm{ani}}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widetilde{\nu}_{\mathcal{E}}}$${\widetilde{\operatorname{\mathsf{A}}}_{G}^{{\mathrm{ani}}}.}$
Conversely, if we assume the existence of a $k$-rational point $\infty_{\rho}$ in the fiber of $\rho$ at $\infty$, we get for any anisotropic $G$-Higgs bundle $F$ an inclusion $\operatorname{\mathsf{Aut}}(F)\to{\mathbb{T}}$ [23, Construction 4.36]. Thus any $k$-point in $I_{\hat{\mu}}\widetilde{{\mathbb{M}}}_{G}^{{\mathrm{ani}}}$ defines a morphism $\kappa:\hat{\mu}\to{\mathbb{T}}$ and in fact a coendoscopic datum ${\mathcal{E}}=(\kappa,\rho_{\kappa},\rho_{\kappa}\to\rho)$ [23, Proposition 5.12]. We say that a coendoscopic datum $\mathcal{E}$} occurs in $I_{\hat{\mu}}\widetilde{{\mathbb{M}}}_{G}^{{\mathrm{ani}}}(k)$ if it arises from a point in $I_{\hat{\mu}}\widetilde{{\mathbb{M}}}_{G}^{{\mathrm{ani}}}(k)$ via this construction and write $I_{\hat{\mu}}\widetilde{{\mathbb{M}}}_{G}^{{\mathrm{ani}}}(k)_{\mathcal{E}}$ for the component with given occurring endoscopy datum $\mathcal{E}$}.
The above construction defines an equivalence [22, Corollary 2.27]
(3.6.2) $\widetilde{\mu}_{\mathcal{E}}:\widetilde{{\mathbb{M}}}_{{\mathcal{E}}}^{G-\infty,{\mathrm{ani}}}(k)\xrightarrow{\>\sim\>}I_{\hat{\mu}}\widetilde{{\mathbb{M}}}_{G}^{{\mathrm{ani}}}(k)_{\mathcal{E}}.$
@@
For our applications we need a variant of this equivalence for the inertia stack of the rigidification $I_{\hat{\mu}}\widetilde{{\mathcal{M}}}_{G}^{{\mathrm{ani}}}$, which will have fewer components indexed by orbits $\mathcal{E}$}] of occuring coendoscopy data under a natural $Z(X,G)$-action [23, Construction 5.20]. Since the coendoscopic groups for two coendoscopy data in the same orbit are isomorphic we will ignore this subtlety and write $I_{\hat{\mu}}\widetilde{{\mathcal{M}}}_{G}^{{\mathrm{ani}}}(k)_{\mathcal{E}}$ for the corresponding component. If we write $\widetilde{{\mathcal{M}}}_{{\mathcal{E}}}^{G}$ for the $Z(X,G)$-rigidification of $\widetilde{{\mathbb{M}}}_{{\mathcal{E}}}^{G-\infty,{\mathrm{ani}}}$, we deduce from (3.6.2) the equivalence [22, Corollary 2.29]
(3.6.3) $\overline{\mu}_{{\mathcal{E}}}:\widetilde{{\mathcal{M}}}_{{\mathcal{E}}}^{G}(k)\xrightarrow{\>\sim\>}I_{\hat{\mu}}\widetilde{{\mathcal{M}}}_{G}^{{\mathrm{ani}}}(k)_{\mathcal{E}}.$
@@
Notice also that because of (3.6.1) the equivalences (3.6.2) and (3.6.3) are compatible with the restriction to individual Hitchin fibers.